OpenStudy (jokersmart):

1 5/9 divide 3 9/25 ? Better view of it below...

OpenStudy (jokersmart):

\[1 \frac{ 5 }{ 9 } \div 3 \frac{ 9 }{ 25 }\] Ah, here we go.

OpenStudy (anonymous):

Can you change that to improper fraction? :)

OpenStudy (anonymous):

in decimal its 1.555555555/3.36

OpenStudy (anonymous):

the answer is 0.46296296294 then im stuck from there

OpenStudy (jokersmart):

Ok, I'll start working on it and I'll show you my work and see if I get anything wrong! Btw, mathway, love the icon! :D Are you part of the bro army? x3

OpenStudy (anonymous):

Sure! Take your time. :)

OpenStudy (anonymous):

bro army? i think no

OpenStudy (anonymous):

What is bro army? O_o

OpenStudy (anonymous):

Once you changed the mixed numbers to improper fractions, divide them. You should be getting a kinda large number, but it can be simplified.

OpenStudy (jokersmart):

\[\frac{ 14 }{ 9 } \times \frac{ 25 }{ 84 }\] This is what I got so far.

OpenStudy (jokersmart):

It's just part of a Youtube thing. If you watched him, you would know haha.

OpenStudy (anonymous):

Now multiply both numerator and denominators. :)

OpenStudy (jokersmart):

Ok! ^_^

OpenStudy (anonymous):

Does it have something to do with Pewdiepie?

OpenStudy (jokersmart):

Yep, yep!

OpenStudy (jokersmart):

Your icon reminded me of it, haha.

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

oh he already gave you the answer... nice

OpenStudy (anonymous):

I don't really like him because his voice annoys me, but of course that's just my opinion. :D

OpenStudy (jokersmart):

I didn't notice that till now. Haha, really? O-o I prefer him than others.

OpenStudy (anonymous):

lml

OpenStudy (jokersmart):

So just 25/54? :>

OpenStudy (anonymous):

Yes. I hope you did your work, and didn't just look at the answer he gave to you. :D

OpenStudy (jokersmart):

I didn't notice his answer till I finished mine. xD I was too busy working on it, but eh.